OpenStudy (anonymous):

(a). If the position of an ant traveling along a horizontal path at time t is 3t^2+1. what is the ant's average velocity from t=1 to t=6 ?

geerky42 (geerky42):

\[\Large v_{avg}=\dfrac{v_f+v_i}{2}\]

OpenStudy (anonymous):

Now we derive ? \[3t ^{2}+1\]

geerky42 (geerky42):

Yes

OpenStudy (anonymous):

so when we derive 3t^2+1 \[=6t\]

geerky42 (geerky42):

yep, that's right.

OpenStudy (anonymous):

so what now?

geerky42 (geerky42):

Did you see the formula I gave you above?

geerky42 (geerky42):

you would need to use that

OpenStudy (anonymous):

yup..!

geerky42 (geerky42):

Can you apply that? vi is velocity at t=1 and vf is velocity at t=6

OpenStudy (anonymous):

\[=\frac{ 6+1 }{ 2 }\] like that?

geerky42 (geerky42):

\(\dfrac{6(6)+6(1)}{2}\)

geerky42 (geerky42):

Because v = 6t

OpenStudy (anonymous):

aha yup ..

OpenStudy (anonymous):

so the final solution .. \[=\frac{ 6(6)+6(1) }{ 2 } = \frac{ 42 }{ 2 } = 21\]

geerky42 (geerky42):

Yep

OpenStudy (anonymous):

hhhhhhhh

OpenStudy (anonymous):

thank you a lot

geerky42 (geerky42):

no problem